Can someone help me on this problem?
What is the perimeter of a rectangle if the length is 5/9 in. and the width is 2/9 in?
Thanks,
sher

Answer by fractalier(6550)

Perimeter P = 2L + 2W, so we have
P = 2(5/9) + 2(2/9)
P = 10/9 + 4/9
P = 14/9
P = 1 5/9 inches